# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

### v1

@@ 2.1. Data Processing for Foundation Models
Foundation models, especially those beyond text, demands innovative data processing paradigms to address the complexity and diversity of large-scale training datasets (Qin et al., 2024b; Zhao et al., 2023; Wang et al., 2023b; Bai et al., 2024b). Despite existing systems for big data or text-only models (Zaharia et al., 2010; Computer, 2023; Soldaini, Luca and Lo, Kyle and Kinney, Rodney and Naik, Aakanksha and Ravichander, Abhilasha and Bhagia, Akshita and Groeneveld, Dirk and Schwenk, Dustin and Magnusson, Ian and Chandu, Khyathi, 2023; Jennings et al., [n.d.]), there is a notable lack of open-source frameworks customized for foundation models due to unique design and implementation challenges:
@@
Functionality Challenges from Data Types and Semantics. Data processing frameworks for foundation models primarily target semi-structured textual and multimodal datasets (e.g., JSONL, Parquet, MP4), often disregarding the strongly structured data types prominent in previous big data systems (e.g., database tables) (Lhoest et al., 2021; Jennings et al., [n.d.]). Significantly, the performance of foundation models depends on the quality, diversity, utility, and safety of data, necessitating a profound semantic understanding and cross-modal alignment to transform raw data into high-quality inputs, thereby enhancing model efficacy (Liu et al., 2023; Li et al., 2023b; Tong et al., 2024).
@@
Efficiency Challenges from Data Operations and Pipelines. Conventional engines support a broad range of SQL queries for complex data operations (Zaharia et al., 2010; Team, 2024a). However, in the context of foundation models, the requirement for traditional grouping and aggregation operations is relatively minimal (Bai et al., 2024b). Instead, emphasis is placed on simpler operations such as Mapper and Filter, which rely heavily on deep learning models via local inference (Schuhmann et al., 2021; Byeon et al., 2022; Wang et al., 2023a) or API calls (Liu et al., 2023; Mei et al., 2024; Maaz et al., 2023). Furthermore, individual records in these datasets can be exceedingly large, surpassing the capabilities of traditional big data engines (Naeem et al., 2022; Ikegwu et al., 2022). When combining multiple data operations into pipelines, it is crucial to maintain efficient data and transformation flows throughout the data lifecycle to prevent bottlenecks stemming from computational overhead, latency, and data integrity.
@@
Usage Challenges from User Interfaces. Traditional big data systems typically use programming interfaces based on Java and Scala, supplemented by SQL. However, practitioners focusing on foundation models predominantly operate within the Python ecosystem, supporting both upstream and downstream repositories and workflows (Paszke et al., 2019; Abadi et al., 2016; Rasley et al., 2020). This shift necessitates designing user interfaces that specifically cater to the new programming needs of foundation model practitioners.
2.2. System Design Goals in Data-Juicer 2.0
@@
Built upon Data-Juicer 1.0, we aim to address challenges identified in Sec. 2.1 with the focus on the following new objectives:
@@
Comprehensive Multimodal Processing Abilities. The first basic goal is to support extensive multimodal processing capabilities across diverse domains, such as AI-generated content, spatial intelligence, education, and entertainment. This involves meticulous processes such as collection, cleansing, annotation, and synthesis, which integrate perceptual and cognitive information across modalities such as video, image, text, and audio (Brooks et al., 2024; Podell et al., 2023; Achiam et al., 2023; Copet et al., 2023).
@@
Efficient Data Processing Operators and Data Flow. Comprehensive semantic operations are expected through advanced procedures that may be high-cost, such as deep learning model invocations, fuzzy deduplication, and clustering. For basic, high-frequency operations, we need to employ dedicated code optimization and speedup techniques like GPU computation or Cython (Behnel et al., 2011). Additionally, optimizing the loading, parsing, and storage of multimodal data, while minimizing overhead associated with linking, read/write operations, and network transmissions, is crucial for streamlining processing and enhancing efficiency.
@@
3. Processing Capabilities Beyond Text
3.1. Overview of New Operators
Data-Juicer 2.0 extends its predecessor, Data-Juicer 1.0—which included approximately 50 text-only OPs—by incorporating over 80 new OPs crafted for multimodal data processing with more functionalities, encompassing modalities such as image, video, and audio. A detailed list of these OPs is available in Table LABEL:tab:mm_ops in the Appendix. For clarity and ease of reference, the OPs are categorized from the following aspects and visually represented in Fig. 2.
  • Modality Types: The majority of new OPs are concentrated on video/image/text-only processing, with approximately 20 OPs dedicated to cross-modal data processing.
  • Operator Types: About 90% are Mappers and Filters, reflecting the user needs in the foundation model domain. New variants in other types, such as Formatters and Deduplicators remain stable due to the broad applicability of already supported formats (e.g., JSONL, Parquet, MP4) and classical algorithms like MinHash (Broder, 1997).
  • Function Types: Data cleaning and analysis operations constitute about half of the new OP suite. Notably, Data-Juicer 2.0 introduces about 40 OPs for data synthesis for cross-modal and post-tuning scenarios, enabling any-to-any generation among 4 supported modalities, and enhancing textual instruction and response based on varied information and needs. Furthermore, new privacy protection OPs and several wrapper OPs for established tools, such as FFmpeg (Developers, [n.d.]), have been integrated, enabling users’ convenient invocation of traditional professional commands.
  • Implementation Types: Model-based OPs dominate, since semantic-aware processing often requires advanced models, such as large models for captioning and visual analysis via SDXL (Podell et al., 2023), GPT (Achiam et al., 2023), and Qwen (Team, 2024c).
(a) Modality type.
(b) Operator type.
(c) Function type.
(d) Implementation type.
Figure 2. Distribution of new OPs across various dimensions.
3.2. Showcase of Typical New Operators
@@
The new OPs include some unique contributions to Data-Juicer 2.0, and others partly inspired by state-of-the-art data processing methodologies for foundation models (Schuhmann et al., 2021; Zhu et al., 2023; Gadre et al., 2024). Below are three representative examples that showcase the diverse computational operations and requirements of these OPs.
@@
phrase_grounding_recall_filter: This OP, newly developed by Data-Juicer 2.0, assesses alignment and consistency between images and textual descriptions. As illustrated in Fig. 3a, it identifies noun phrases in the text that refer to key entities. Subsequently, an open-vocabulary object detection model, such as Owl-ViT (Minderer et al., 2022), attempts to detect corresponding entities within the image. The OP then calculates the recall of detected phrases to evaluate consistency, where a higher recall indicates better image-text coherence.
@@
video_motion_score_filter: This OP quantifies video dynamics. As shown in Fig. 3b, it samples multiple frames at a specified frames-per-second (FPS) rate and computes optical flows. The average magnitude of these flows determines the motion score, with higher scores indicating more dynamic content. To accommodate various computational resources, both a GPU-based RAFT version (Teed and Deng, 2020) and a CPU-based OpenCV version (Bradski, 2000) are available.
@@
Figure 3. Illustrative examples of new operators.
4. User Interfaces
4.1. Low-Level APIs
@@
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental DJ-Dataset, DJ-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded DJ-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated DJ-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the DJ-Dataset and DJ-Operators will be introduced later in Sec. 5. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
4.2. RESTful APIs
@@
Utilizing standard Python type hints, we provide one-click generation of high-performance web APIs, capable of automatically discovering, registering, and adapting OP classes and related tools. Users can rapidly initiate a web server supporting the Asynchronous Server Gateway Interface by simply executing a service script, eliminating the need for manual code writing. The asynchronous concurrency mechanism enables options such as lightweight background tasks and mitigates potential bottlenecks for endpoints that may experience prolonged network I/O blocking. Each OP is accessible via POST requests, typically executing the OP’s run() method as the endpoint. The target dataset path is passed through query parameters, with additional configurable OP parameters transmitted via JSON payload. Upon completion, the path to the processed dataset is returned. This invocation through Web APIs allows for the centralized host with distributed access, reducing deployment complexity. It also facilitates the separation of application logic from execution, potentially fostering the development and release of more applications built upon Data-Juicer. Importantly, the extensive customization parameters available in the programming API can be seamlessly passed through the Web API, maintaining full functionality without compromising usability, and facilitating serviceful calling by higher-level interfaces as introduced later.
4.3. Web & CMD Tools
Figure 4. An example of config file for data recipe.

### v2

@@ A.1 Data Processing for Foundation Models
Foundation models, especially those beyond text, demands innovative data processing paradigms to address the complexity and diversity of large-scale training datasets [54, 79, 74, 5]. Despite existing systems for big data or text-only models [78, 21, 58, 32], there is a notable lack of open-source frameworks customized for foundation models due to unique design and implementation challenges:
@@
Functionality Challenges from Data Types and Semantics. Data processing frameworks for foundation models primarily target semi-structured textual and multimodal datasets (e.g., JSONL, Parquet, MP4), often disregarding the strongly structured data types prominent in previous big data systems (e.g., database tables) [36, 32]. Significantly, the performance of foundation models depends on the quality, diversity, utility, and safety of data, necessitating a profound semantic understanding and cross-modal alignment to transform raw data into high-quality inputs, thereby enhancing model efficacy [40, 38, 70].
@@
Efficiency Challenges from Data Operations and Pipelines. Conventional engines support a broad range of SQL queries for complex data operations [78, 60]. However, in the context of foundation models, the requirement for traditional grouping and aggregation operations is relatively minimal [5]. Instead, emphasis is placed on simpler operations such as Mapper and Filter, which rely heavily on deep learning models via local inference [57, 11, 73] or API calls [40, 43, 41]. Furthermore, individual records in these datasets can be exceedingly large, surpassing the capabilities of traditional big data engines [47, 31]. When combining multiple data operations into pipelines, it is crucial to maintain efficient data and transformation flows throughout the data lifecycle to prevent bottlenecks stemming from computational overhead, latency, and data integrity.
@@
Usage Challenges from User Interfaces. Traditional big data systems typically use programming interfaces based on Java and Scala, supplemented by SQL. However, practitioners focusing on foundation models predominantly operate within the Python ecosystem, supporting both upstream and downstream repositories and workflows [49, 1, 55]. This shift necessitates designing user interfaces that specifically cater to the new programming needs of foundation model practitioners.
A.2 System Design Goals in Data-Juicer 2.0
@@
Built upon Data-Juicer 1.0, we aim to address challenges identified in Appendix A.1 with the focus on the following new objectives:
@@
Comprehensive Multimodal Processing Abilities. The first basic goal is to support extensive multimodal processing capabilities across diverse domains, such as AI-generated content, spatial intelligence, education, and entertainment. This involves meticulous processes such as collection, cleansing, annotation, and synthesis, which integrate perceptual and cognitive information across modalities such as video, image, text, and audio [10, 52, 2, 22].
@@
Efficient Data Processing Operators and Data Flow. Comprehensive semantic operations are expected through advanced procedures that may be high-cost, such as deep learning model invocations, fuzzy deduplication, and clustering. For basic, high-frequency operations, we need to employ dedicated code optimization and speedup techniques like GPU computation or Cython [6]. Additionally, optimizing the loading, parsing, and storage of multimodal data, while minimizing overhead associated with linking, read/write operations, and network transmissions, is crucial for streamlining processing and enhancing efficiency.
@@
These design goals guide the development of robust operator functionalities and system implementation, tailored to meet the specific requirements of foundation model data processing. In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
Appendix B List of New Operators in Data-Juicer 2.0
The full list of new OPs in Data-Juicer 2.0 is shown in Table LABEL:tab:new_ops.
{longtblr}
[ caption = List of new OPs in Data-Juicer 2.0. The full OP list can be found in our document in GitHub repository., label = tab:new_ops, entry = none, ] width = colspec = Q[30]Q[200]Q[400]Q[85]Q[80], cell21 = r=12, cell141 = r=12, cell261 = r=27, cell531 = r=8, cell611 = r=5, cell661 = r=25, Modality Type OP Name (OP Type as the last suffix) OP Description Function Type Implementation Type
Image-only image_blur_mapper Blur images Augmentation Rule-based
image_remove_background_mapper Remove background of images Augmentation Model-based
image_face_blur_mapper Blur faces detected in images Privacy Model-based
image_aesthetics_filter Keeps samples containing images whose aesthetics scores are within the specified range Cleaning Model-based
image_aspect_ratio_filter Keeps samples containing images with aspect ratios within the specified range Cleaning Rule-based
image_face_ratio_filter Keeps samples containing images with face area ratios within the specified range Cleaning Model-based
image_nsfw_filter Keeps samples containing images with NSFW scores below the threshold Cleaning Model-based
image_shape_filter Keeps samples containing images with widths and heights within the specified range Cleaning Rule-based
image_size_filter Keeps samples containing images whose size in bytes are within the specified range Cleaning Rule-based
image_watermark_filter Keeps samples containing images with predicted watermark probabilities below the threshold Cleaning Model-based
image_deduplicator Deduplicates samples at document-level using exact matching of images between documents Deduplication Model-based
ray_image_deduplicator Deduplicates samples at document-level using exact matching of images between documents on ray Deduplication Model-based
Image-Text image_captioning_from_gpt4v_mapper Generate texts based on GPT-4V for images Synthesis Model-based
image_captioning_mapper Generate texts based on image-to-text models for images Synthesis Model-based
image_diffusion_mapper Generate images based on text-to-image models for texts Synthesis Model-based
mllm_mapper Use multimodal large language models for visual question answering tasks Wrapper of external tool  Model-based
sdxl_prompt2prompt_mapper Use the generative model SDXL and image editing technique Prompt-to-Prompt to generate pairs of similar images Synthesis Model-based
image_segment_mapper Perform segment-anything on images and return the bounding box values Wrapper of external tool  Model-based
sentence_augmentation_mapper Augment sentences using large language models Augmentation Model-based
image_text_matching_filter Keeps samples with image-text classification matching score within the specified range based on a BLIP model Cleaning Model-based
image_text_similarity_filter Keeps samples with image-text feature cosine similarity within the specified range based on a CLIP model Cleaning Model-based
phrase_grounding_recall_filter Keeps samples whose locating recalls of phrases extracted from text in the images are within a specified range Cleaning Model-based
image_pair_similarity_filter Keeps image pairs whose cosine similarity of features is within a specified range, based on a CLIP model. Cleaning Model-based
text_pair_similarity_filter Keeps text pairs whose cosine similarity of features is within a specified range, based on a CLIP model Cleaning Model-based
Video-only video_face_blur_mapper Blur faces detected in videos Privacy Model-based
video_ffmpeg_wrapped_mapper Wrapper to run a FFmpeg video filter Wrapper Tool-based
video_remove_watermark_mapper Remove watermarks in videos Cleaning Model-based
video_resize_aspect_ratio_mapper Resize video aspect ratio to a specified range Augmentation Rule-based
video_resize_resolution_mapper Map videos to ones with given resolution range Augmentation Rule-based
video_split_by_duration_mapper Split video by duration Augmentation Rule-based
video_spit_by_key_frame_mapper Split video by key frames Augmentation Rule-based
video_split_by_scene_mapper Split videos into scene clips Augmentation Model-based
video_human_tracks_extraction_mapper Extracts human tracks by linking face and body bounding boxes across frames Synthesis Model-based
video_human_demographics_mapper determines demographic attributes (gender, age, race) for face tracks by aggregating frame-level detections. Synthesis Model-based
video_human_description_mapper generates individual-focused videos from body bounding box tracks and processes it for appearance description and simple actions. Synthesis Model-based
video_facial_description_mapper generates face-focused videos from face bounding box tracks and processes them for emotion and facial description. Synthesis Model-based
video_active_speaker_detection_mapper detects active speaking by analyzing face track sequences alongside corresponding audio Synthesis Model-based
video_ASR_mapper generate the automatic speech recognition result for video Synthesis Model-based
video_speech_emotion_recognition_mapper Detect the emotion category of the speech in the video Synthesis Model-based
video_voice_demographics_mapper analyzes voice demographics (such as gender, age) in video Synthesis Model-based
video_aesthetics_filter Keeps samples whose specified frames have aesthetics scores within the specified range Cleaning Model-based
video_aspect_ratio_filter Keeps samples containing videos with aspect ratios within the specified range Cleaning Rule-based
video_duration_filter Keep data samples whose videos’ durations are within a specified range Cleaning Rule-based
video_face_ratio_filter Keep samples whose frame ratio containing faces is greater than a certain threshold are retained. Cleaning Model-based
video_motion_score_filter Keep samples with video motion scores within a specific range Cleaning Rule-based
video_nsfw_filter Keeps samples containing videos with NSFW scores below the threshold Cleaning Model-based
video_ocr_area_ratio_filter Keep data samples whose detected text area ratios for specified frames in the video are within a specified range Cleaning Model-based
video_resolution_filter Keeps samples containing videos with horizontal and vertical resolutions within the specified range Cleaning Rule-based
video_watermark_filter Keeps samples containing videos with predicted watermark probabilities below the threshold Cleaning Model-based
video_deduplicator Deduplicates samples at document-level using exact matching of videos between documents Deduplication Model-based
ray_video_deduplicator Deduplicates samples at document-level using exact matching of videos between documents on ray Deduplication Model-based
Video-Text video_captioning_from_audio_mapper Generate texts for videos according to their audio streams based on audio LLMs Synthesis Model-based
video_captioning_from_frames_mapper Generate texts for videos according to sampled frame images based on image-to-text models Synthesis Model-based
video_captioning_from_video_mapper Generate texts for videos based on video-to-text models Synthesis Model-based
video_tagging_from_audio_mapper Generate tags from audio streams extracted from the video. Synthesis Model-based
video_tagging_from_frames_mapper Generate video tags from frames extracted from the video. Synthesis Model-based
video_captioning_from_summarizer_mapper Generate texts by summarizing several types of generated texts (from video/audio/frames, tags from audio/frames, …) Synthesis Model-based
video_frames_text_similarity_filter Keep data samples whose similarities between sampled video frame images and text are within a specific range Cleaning Model-based
video_tagging_from_frames_filter Keep samples containing videos with given tags Cleaning Model-based
Audio-only audio_ffmpeg_wrapped_mapper Wrapper to run a FFmpeg audio filter Wrapper Tool-based
audio_add_gaussian_noise_mapper Add gaussian noise to audio. Augmentation Tool-based
audio_duration_filter Keep data samples whose audios’ durations are within a specified range Cleaning Rule-based
audio_nmf_snr_filter Keep data samples whose audios’ Signal-to-Noise Ratios are within a specified range Cleaning Rule-based
audio_size_filter Keep data samples whose audios’ sizes are within a specified range Cleaning Rule-based
Text-only calibrate_qa_mapper Calibrate question-answer pairs based on reference text Augmentation Model-based
calibrate_query_mapper Calibrate query in question-answer pairs based on reference text Augmentation Model-based
calibrate_response_mapper Calibrate response in question-answer pairs based on reference text Augmentation Model-based
extract_entity_attribute_mapper Extract attributes for given entities from the text. Synthesis Model-based
extract_entity_relation_mapper Extract entities and relations in the text for knowledge graph. Synthesis Model-based
extract_event_mapper Extract events and relevant characters in the text. Synthesis Model-based
extract_keyword_mapper Generate keywords for the text. Synthesis Model-based
extract_nickname_mapper Extract nickname relationship in the text. Synthesis Model-based
extract_support_text_mapper Extract support sub text for a summary. Synthesis Model-based
extract_tables_from_html_mapper Extract tables from HTML content. Cleaning Rule-based
generate_qa_from_examples_mapper Generate question and answer pairs based on examples. Synthesis Model-based
generate_qa_from_text_mapper Generate question and answer pairs from text. Synthesis Model-based
optimize_qa_mapper Optimize both the query and response in question-answering samples. Augmentation Model-based
optimize_query_mapper Optimize the query in question-answering samples. Augmentation Model-based
optimize_response_mapper Optimize the response in question-answering samples. Augmentation Model-based
pair_preference_mapper Construct paired preference samples. Augmentation Model-based
text_chunk_mapper Split input text to chunks. Synthesis Model-based
naive_grouper Group all samples to one batched sample. Synthesis Rule-based
key_value_grouper Group samples to batched samples according values in given keys. Synthesis Rule-based
entity_attribute_aggregator Return conclusion of the given entity’s attribute from some docs. Synthesis Model-based
most_relavant_entities_aggregator Return most relevant entities with the given entity from some docs. Synthesis Model-based
nested_aggregator Considering the limitation of input length, nested aggregate contents for each given number of samples. Synthesis Model-based
llm_quality_score_filter Keep sample with high quality score estimated by LLM. Cleaning Model-based
llm_difficulty_score_filter Keep sample with high difficulty score estimated by LLM. Cleaning Model-based
domain_diversity_selector Select samples based on the data’s domain diversity. Cleaning Model-based
Appendix C Showcase of Typical Multimodal Operators
@@
The new OPs include some unique contributions to Data-Juicer 2.0, and others partly inspired by SOTA data processing methodologies for foundation models [57, 83, 26]. Below are three representative examples that showcase the diverse computational operations and requirements of these OPs.
@@
phrase_grounding_recall_filter: This OP, newly developed by Data-Juicer 2.0, assesses alignment and consistency between images and textual descriptions. As illustrated in Fig. 5a, it identifies noun phrases in the text that refer to key entities. Subsequently, an open-vocabulary object detection model, such as Owl-ViT [44], attempts to detect corresponding entities within the image. The OP then calculates the recall of detected phrases to evaluate consistency, where a higher recall indicates better image-text coherence.
@@
video_motion_score_filter: This OP quantifies video dynamics. As shown in Fig. 5b, it samples multiple frames at a specified frames-per-second (FPS) rate and computes optical flows. The average magnitude of these flows determines the motion score, with higher scores indicating more dynamic content. To accommodate various computational resources, both a GPU-based RAFT version [68] and a CPU-based OpenCV version [7] are available.
@@
Figure 5: Illustrative examples of new operators.
Appendix D User Interfaces
D.1 Low-level APIs
@@
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental Data-Juicer-Dataset, Data-Juicer-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded Data-Juicer-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated Data-Juicer-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the Data-Juicer-Dataset and Data-Juicer-Operators are introduced in Sec. 4 and Appendix E. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
D.2 RESTful APIs
@@
Utilizing standard Python type hints, we provide one-click generation of high-performance web APIs, capable of automatically discovering, registering, and adapting OP classes and related tools. Users can rapidly initiate a web server supporting the Asynchronous Server Gateway Interface by simply executing a service script, eliminating the need for manual code writing. The asynchronous concurrency mechanism enables options such as lightweight background tasks and mitigates potential bottlenecks for endpoints that may experience prolonged network I/O blocking. Each OP is accessible via POST requests, typically executing the OP’s run() method as the endpoint. The target dataset path is passed through query parameters, with additional configurable OP parameters transmitted via JSON payload. Upon completion, the path to the processed dataset is returned. This invocation through Web APIs allows for a centralized host with distributed access, reducing deployment complexity. It also facilitates the separation of application logic from execution, potentially fostering the development and release of more applications built upon Data-Juicer. Importantly, the extensive customization parameters available in the programming API can be seamlessly passed through the Web API, maintaining full functionality without compromising usability, and facilitating serviceful calling by higher-level interfaces as introduced later.
D.3 Web & CMD tools
Figure 6: An example of config file for data recipe.